# Learning the Intrinsic Dimensionality of Fermi-Pasta-Ulam-Tsingou Trajectories: A Nonlinear Approach using a Deep Autoencoder Model

### v1

@@ Abstract
We address the intrinsic dimensionality (ID) of high-dimensional trajectories, comprising $n_{s}=4\,000\,000$ data points, of the Fermi-Pasta-Ulam-Tsingou (FPUT) $\beta$ model with $N=32$ oscillators. To this end, a deep autoencoder (DAE) model is employed to infer the ID in the weakly nonlinear regime ($\beta\lesssim 1$). We find that the trajectories lie on a nonlinear manifold of dimension $m^{\ast}=2$ embedded in a $64$-dimensional phase space. The DAE further reveals that this dimensionality increases to $m^{\ast}=3$ at $\beta=1.1$, coinciding with a symmetry breaking transition, in which additional energy modes with even wave numbers $k=2,4$ become excited. Finally, we discuss the limitations of the linear approach based on principal component analysis (PCA), which fails to capture the underlying structure of the data and therefore yields unreliable results in most cases.
@@ I Introduction
Fermi believed that weakly coupled nonlinear systems would inevitably exhibit ergodic behavior, as required for approaching thermal equilibrium [20]. As a consequence, in 1953, the first computer simulations of the Fermi-Pasta-Ulam-Tsingou (FPUT) model, a one-dimensional chain of weakly nonlinearly coupled oscillators, started aiming to study the system’s thermalization, which in turn should have confirmed the validity of ergodic hypothesis [16, 21, 74, 23, 44, 14]. In contrast, the energy recurrence phenomenon observed from simulations, for which the total energy was shared between a few energy modes only, seemed to contradict the ergodic hypothesis, a fact often referred to as the FPUT paradox.
@@
Since then, the FPUT model has had a huge impact on our understanding of nonlinear dynamics and deterministic chaos. It would be an impossible task to mention all the important contributions made in the last seventy years, attempting to explain the peculiar dynamical features of FPUT chain. Here, we limit ourselves to recall that it was found that the recurrence phenomenon can be linked to Kolmogorov-Arnol’d-Moser (KAM) theorem [42, 3, 55], Toda lattice  [1, 5], the solitonic behaviour of Korteweg de Vries (KdV) equation [76] and q-breathers [17, 18]. On the other hand, there exist various theories and methods attempting to explain how eventually the system approaches the equilibrium: existence of a stochastic threshold [35, 48], the FPUT metastability [22, 23, 65], and the spectral-entropy equipartition indicator [47, 23].
@@
Recently, a data-driven approach based on principal component analysis (PCA) was proposed to infer the intrinsic (or effective) dimensionality (ID) $m^{\ast}$ of the high-dimensional trajectories of the FPUT $\beta$ model with $N=32$ oscillators [52]. This work was chiefly motivated by the observation that the ergodic hypothesis requires trajectories to lie on a constant-energy hypersurface of dimension n1n-1 in $\mathbb{R}^{n}$ (n=2Nn=2N[34]. As a result, thermal equilibrium cannot be reached when the trajectories lie on very low-dimensional manifolds, a scenario which aligns with the application of KAM theorem to FPUT model, in a weakly nonlinear regime [4]. In fact, accordingly to KAM theorem the orbits should be constrainted to low-dimensional invariant tori [19, 66]. In this regard, PCA when applied to entire trajectory data consisting of $n_{s}=4\,000\,000$ data points, with the initial condition corresponding to excitation of only the first energy mode ($k=1$), estimates the ID to be either $2$ or $3$ when $\beta\lesssim 1.1$. These findings strongly suggest a non-ergodic behavior, and are consistent with other qualitative observations that support the underlying quasi-periodic motion of FPUT system: the hexagonal patterns shown by Poincaré maps [26], and the energy recurrences observed from the simulations [52].
@@
However, being linear, PCA provides a rather limited approach to the problem at hand [37, 30, 45, 28, 54, 69]. Indeed, it assumes that the data points lie on optimal hyperplanes of a given rank $m$ (or equivalently, on $m$-dimensional linear subspaces if data is centered), which is not necessarily the case, as shown by a qualitative analysis of early-stage orbits, performed using $t$-distributed stochastic neighbor embedding ($t$-SNE), a nonlinear manifold learning algorithm [73, 40, 41, 12, 52]. More importantly, the ineherent limitation of a linear approach was further confirmed by the analysis of trajectory data for $\beta=0.1$, provided by the multi-chart flows, a Riemannian manifold learning method [75]. Specifically, this state-of-the-art method suggests that the trajectory lies near a nonlinear Riemannian manifold of dimension $m^{\ast}=2$. Curiously, in such a case Participation Ratio (PR) heuristic [43] (Eq. 10), applied to the PCA eigenvalues, yields the very same result, but only accidentally: in fact, PCA would be expected to overestimate the intrinsic dimensionality by one, thus predicting $m^{\ast}=3$ [77] (see Sect. IV).
@@
Accordingly, in this work, we shall replace the PCA with its nonlinear generalization: the autoencoder (AE) [27, 62]. The AE is a type of artificial neural network (ANN), capable of learning the nonlinear structure of the data. In particular, we shall employ a relatively simple undercomplete deep autoencoder (DAE) model for learning the nonlinear structure of trajectory data (see Sect. III). We found that this model can learn and generalizes appropriately, yielding negligible reconstruction errors in comparison to those obtained using PCA. From the DAE reconstruction error curves, the corresponding elbow points estimate $m^{\ast}=2$ for all available trajectory data associated with the recurrence phenomenon $\beta\lesssim 1$ (see Fig. 4).
@@
Furthermore, the DAE model predicts that the intrinsic dimensionality increases by unity, i.e., $m^{\ast}=3$, when $\beta=1.1$. We find that this change is associated with the onset of a symmetry breaking (SB) phenomenon characteristic of the $\beta$ model, whereby additional energy modes with even wave numbers, $k=2,4$, begin to share the system’s total energy. In contrast to the DAE model, this important variation in the ID is completely missed by principal component analysis, for which the estimated ID remains constant as $\beta\rightarrow 1.1$, regardless of the suitable heuristic employed, such as Kneedle algorithm (KA) [68], PR or Kaiser criterion (KC) [36], to estimate it.
@@ II FPUT $\beta$ Model
Accordingly, the $\beta$ model’s equations of motion are the following  [^1] The equations of motion are commonly referred to as the canonical Hamilton equations.
@@
From a mathematical point of view, the Cauchy problem for the first-order system in Eq. 2 is well posed; that is, the solution exists and is unique because the vector field is analytic, being polynomial, as the number of oscillators $N$ is finite [53]. However, the solution can be obtained only through numerical integration. To this end, we employed the velocity Störmer-Verlet scheme, which is a second-order symplectic integrator well suited for Hamiltonian systems like FPUT chains [29, 6]. Accordingly, in the present work, the trajectories were generated with high accuracy using a fixed time step of 0.05 [26].
@@ III A Deep Autoencoder Model
The autoencoder is an ANN trained to reconstruct an input vector $\mathbf{x}\in\mathbb{R}^{n}$ at the output layer, i.e., to produce a reconstruction $\mathbf{\tilde{x}}$, by passing $\mathbf{x}$ through a narrower hidden layer referred to as the latent (or bottleneck) layer. In this layer, the code vector $\mathbf{z}\in\mathbb{R}^{m}$ ($1\leq m\ll n$) provides a compressed, low-dimensional, and salient representation of $\mathbf{x}$. This representation is achieved by minimizing the reconstruction error $J_{m}$ (see Eq.  8), or equivalently, the mean squared error (MSE) loss. In general, AEs can be regarded as nonlinear generalizations of PCA: indeed, it can be shown that an autoencoder restricted to be linear is equivalent to PCA [46, 64].
In the present work, we shall employ a relatively simple deep autoencoder model with five hidden layers, whose architecture is schematically depicted in the cartoon shown in Fig. 2. The DAE architecture is composed of two sequential networks: the encoder and the decoder. The first maps $\mathbf{x}$ to $\mathbf{z}$, while the second maps $\mathbf{z}$ to $\mathbf{\tilde{x}}$.
Below, we describe all the model’s hyperpameters chosen according to its excellent reconstruction performance of the input data for latent dimension $m\geq 2$ as discussed in Sect. IV. The number of nodes (or neurons) per layer together with the respective activation functions are shown in Table 1. Note that all hidden nodes use the popular ReLU (Rectified Linear Unit) activation function, i.e., $h(y)=\max(0,y)$, for learning nonlinear mappings [57]. By contrast, the output layer employs a linear activation function, which is commonly used for regression tasks. However, it would also be possible to choose a linear activation function for the bottleneck layer, since no significant differences in performance are observed.
Figure 2: Schematic view of a DAE model with five hidden layers. The input and output layers contain $n$ nodes, while the bottleneck layer has $m$ nodes ($m\ll n$). All hidden nodes use ReLU activation function, as indicated by the symbol inside the nodes. The output layer’s nodes employ a linear activation function which is appropriate for regression.
Layer Type Nodes Activation Function
@@
hidden layer $m$ ReLu
@@
Table 1: Number of nodes and respective activation functions of DAE architecture.
@@
The model is trained using the Adam optimizer [39] with a default learning rate of $0.001$ and a mini-batch size of $256$, noting that smaller batch sizes (e.g., $32$) do not produce significant differences in performance. The number of training iterations is set to $300$ epochs. However, early stopping (ES), which acts as an implicit regularizer with a patience of $30$ epochs, is chiefly used to avoid unnecessary training, thereby saving significant time and computational resources. The model is implemented using Keras and TensorFlow [25, 11, 2].
@@
In general, especially for neural networks, the training reconstruction error can be very low because the model may fit the training data extremely closely, thereby leading to serious overfitting. As a result, the model may not perform well on unseen data. Accordingly, to ensure that the model generalizes successfully, we split each trajectory dataset into a training set and a test set [31]. The training set contains $70\%$ of the input data, consisting of data points chosen at random, while the test set contains the remaining $30\%$ of the data. Note that $20\%$ of training set is also used as validation set for monitoring the model’s generalization during training. With this choice, we find that the difference between the training and test reconstruction errors is generally $\lesssim 10^{-3}$, and decreases to about $10^{-5}$ for larger bottleneck-layer sizes, i.e., as $m\rightarrow 7$. Overall, these errors are smaller by at least a factor of $100$ than the corresponding PCA reconstruction errors. This fact is illustrated for the specific case of $\beta=1.1$ in the top panel of Fig. 3, where the DAE and PCA reconstruction curves are compared. Therefore, the DAE model provides a significantly more accurate reconstruction of the input data than PCA.
Here, it is worth noting that when the latent dimensionality is restricted to a single variable ($m=1$), it is unrealistic to expect an autoencoder to capture the complex structure of the $64$-dimensional input data, unless the data lie on an approximately one-dimensional manifold [46]. This condition is not satisfied in the present case, as the intrinsic dimensionality of the trajectory data satisfies $\geq 2$. Consequently, the severe information bottleneck imposed by $m=1$ leads to underfitting and, therefore, to increased reconstruction errors. Nevertheless, there is no need to modify the present model, as the aforementioned limitation disappears altogether for latent dimension $m>1$. Instead, we rely on early stopping to retain the model corresponding to the minimum validation error during the optimization of the trainable parameters (weights and biases)  [^2] Note that when the model’s bottleneck has a single node, i.e., $m=1$, there are $4,801$ trainable parameters.. A detailed analysis of the DAE’s learning curves for $\beta=0.3$ is provided in Sect. IV.
@@ III.1 Datasets and Data Preprocessing
Each dataset of size $n_{s}=4\,000\,000$ corresponds to a set $\mathcal{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{i},\cdots\mathbf{x}_{n_{s}}\}$, whose elements $\mathbf{x}_{i}\in\mathbb{R}^{n}$ are data points belonging to a single trajectory obtained from numerical simulations of the FPUT system at a given value of $\beta\in[0.1,1.1]$, sampled at intervals of $\Delta\beta=0.1$ [52].
The set $\mathcal{X}$ is then conveniently arranged into an $n_{s}\times n$ data matrix $X$. However, before feeding it into the DAE model, a data preprocessing step is required, consisting of centering and scaling each variable by its standard deviation, as is customary in PCA [37]. First, centering the data,corresponding to setting the mean of each variable to zero, allows the PCA eigenvalues to be efficiently computed via singular value decomposition (SVD) (see Sect. B). Second, rescaling ensures that each variable, now having unit variance, contributes equally to the analysis regardless of its original scale [37]. This preprocessing step allows for a direct comparison between the DAE results and those obtained from principal component analysis in Ref. [52].
@@ IV Results and Discussion
To start with, we briefly analyze the learning curves obtained during model training on trajectory data with $\beta=0.5$, which exemplify the typical DAE performance across the datasets considered.
In Fig. 3 (main panel), the training and validation mean squared error losses are shown as functions of epoch for bottleneck sizes $m=1$ and $m=2$. In both cases, the MSE losses exhibit a rapid initial decay followed by plateau, indicating that the model quickly learns the dominant structure of the data before reaching an error floor, that is, the minimum achievable error. Moreover, for each value of $m$, the training and validation curves remain close throughout training, suggesting that the DAE generalizes well to unseen data (no overfitting).
However, when the bottleneck dimension is reduced to $m=1$, the loss saturates at a significantly higher value than for $m=2$. Specifically, considering that early stopping is triggered at around epoch $116$, one finds that the minimum validation MSE loss $\approx 0.036$ which is attained at epoch $86$. This is roughly $7$ times larger than the corresponding MSE loss observed at $m=2$  [^3] The best validation MSE loss is $\approx 0.0049$ attained at epoch $101$ when $m=2$.. Additionally, the model’s capacity limitation is further evidenced by the noisier evolution of the loss curves for $m=1$, including the presence of spikes and late-epoch fluctuations. These facts are indicative of optimization difficulties induced by the scalar bottleneck and cannot be substantially mitigated without increasing the latent dimensionality, regardless of changes to the model architecture or hyperparameters, such as the learning rate. By contrast, for $m\geq 2$ the model performs well, yielding substantially smaller reconstruction errors, as further supported by the well-converged validation loss curves for $m=3,4,5$ shown in the inset of Fig. 3.
@@
Figure 3: Training and validation MSE loss curves as functions of epoch for bottleneck sizes $m=1,2$ and trajectory data with $\beta=0.3$. (Inset) Validation MSE loss curves as functions of epoch for $m=3,4,5$ and same data. Throughout training, the deep autoencoder employs an early stopping with a patience of $30$.
@@
After training the deep autoencoder, the reconstruction errors $J_{m}$ obtained from the trained model for a given value $\beta$ form a curve that depends on the latent dimension $m$. The intrinsic dimensionality can then be inferred from the elbow of the resulting reconstruction error curve [9]. This heuristic rests on the assumption that, beyond the elbow, $J_{m}$ no longer decreases significantly as $m$ increases [72]. Note that this approach is inherently subject to a degree of ambiguity and subjectivity, but still routinely employed in machine learning: manifold reduction and clustering [25]. In the following we shall automate the identification of the elbow using the Kneedle algorithm, even though in the present case the elbow points can be easily identified by visual inspection.
@@
Figure 4: (Top panel) DAE test reconstruction error curves of the test data as functions of the dimension $m$ for $\beta=0.1,0.4,0.7,0.8,0.9,1.0$, and $1.1$. The PCA reconstruction error curve, corresponding to data with $\beta=1.1$, is rescaled by a factor of $100$ for proper visualization. (Bottom panel) DAE reconstruction error curve for the test data corresponding to $\beta=0.1$.

### v2

@@ Abstract
We address the intrinsic dimensionality (ID) of high-dimensional trajectories, comprising $n_{s}=4\,000\,000$ data points, of the Fermi-Pasta-Ulam-Tsingou (FPUT) $\beta$ model with $N=32$ oscillators. To this end, a deep autoencoder (DAE) is used to infer the ID in the weakly nonlinear regime where energy recurrences are observed ($\beta\lesssim 1$). We find that the trajectories lie on a nonlinear Riemannian manifold of dimension $m^{\ast}=2$ embedded in a $64$-dimensional phase space. By contrast, principal component analysis (PCA) together with the Participation Ratio (PR) method provides only a reasonable upper bound on the ID for each value of $\beta$. Our DAE further reveals that the ID increases to $m^{\ast}=3$ at $\beta=1.1$, coinciding with a symmetry-breaking (SB) phenomenon characteristic of the $\beta$ model, in which additional energy modes with even wave numbers $k=2,4$ become excited. Notably, the SB phenomenon cannot be detected by the linear approach provided by PCA.
@@ I Introduction
Fermi believed that weakly coupled nonlinear systems would inevitably exhibit ergodic behavior, as required to approach thermal equilibrium [22]. As a consequence, in 1953, the first computer simulations of the Fermi-Pasta-Ulam-Tsingou (FPUT) model, a one-dimensional chain of weakly nonlinearly coupled oscillators, were started aiming to study the system’s thermalization, which in turn should have confirmed the validity of the ergodic hypothesis [18, 23, 80, 26, 49, 15]. In contrast, the energy recurrence phenomenon observed from simulations, for which the total energy was shared between a few energy modes only, seemed to contradict the ergodic hypothesis, a fact often referred to as the FPUT paradox.
@@
Since then, the FPUT model has had a huge impact on our understanding of nonlinear dynamics and deterministic chaos. It would be an impossible task to mention all the important contributions made in the last seventy years attempting to explain the peculiar dynamical features of the FPUT chain. Here, we limit ourselves to recalling that the recurrence phenomenon can be linked to the Kolmogorov–Arnol’d–Moser (KAM) theorem [47, 3, 60], the Toda lattice [1, 5], the solitonic behaviour of the Korteweg–de Vries (KdV) equation [82], and q-breathers [19, 20]. Additionally, there exist various theories and methods attempting to explain how the system eventually approaches equilibrium: the existence of a stochastic threshold [39, 53], the FPUT metastability [25, 26, 70], and the spectral-entropy equipartition indicator [52, 26].
@@
Recently, a data-driven approach based on principal component analysis (PCA) was proposed to infer the intrinsic (or effective) dimensionality (ID) $m^{\ast}$ of the high-dimensional trajectories of the FPUT $\beta$ model with $N=32$ oscillators [57]. This work was chiefly motivated by the observation that the ergodic hypothesis requires trajectories to lie on a constant-energy hypersurface of dimension n1n-1 in $\mathbb{R}^{n}$ (n=2Nn=2N[38]. As a result, thermal equilibrium cannot be reached when the trajectories lie on very low-dimensional manifolds, a scenario which aligns with the application of the KAM theorem to the FPUT model in a weakly nonlinear regime [4]. Indeed, according to the KAM theorem, the orbits should be constrained to low-dimensional invariant tori [21, 71].
In this regard, PCA when applied to entire trajectory data consisting of $n_{s}=4\,000\,000$ data points, with the initial condition corresponding to excitation of only the first energy mode ($k=1$), estimates the ID to be either $2$ or $3$ when $\beta\lesssim 1.1$ [57]. These findings strongly suggest non-ergodic behavior, and are consistent with other qualitative observations in the same regime that support the underlying quasi-periodic motion of the FPUT system: the hexagonal patterns shown by Poincaré maps [30], and the energy recurrences observed from the simulations [57].
@@
However, being linear, PCA provides a rather limited approach to the problem at hand [42, 34, 50, 32, 59, 74]. Indeed, it assumes that the data points lie on optimal hyperplanes of a given rank $m$ (or, equivalently, on $m$-dimensional linear subspaces if the data are centered). This assumption is not necessarily satisfied, as shown by a qualitative analysis of early-stage orbits performed using $t$-distributed stochastic neighbor embedding ($t$-SNE), a nonlinear manifold-learning algorithm [78, 45, 46, 13, 57]. More importantly, the inherent limitation of a linear approach was further confirmed by the analysis of trajectory data for $\beta=0.1$ using multi-chart flows, a manifold-learning method [81]. Specifically, this state-of-the-art method suggests that the trajectory lies near a nonlinear Riemannian manifold of dimension $m^{\ast}=2$.
@@
With this in mind, it makes sense to replace PCA with its nonlinear generalization: the autoencoder (AE) [31, 67]. The AE is a type of artificial neural network (ANN) capable of learning the nonlinear structure of data. Specifically, we employ a relatively simple undercomplete deep autoencoder (DAE) to learn the actual nonlinear structure of the trajectory data under scrutiny (see Sec. III). We find that this model can learn and generalise appropriately, yielding negligible reconstruction errors in comparison to those obtained using PCA. From the DAE reconstruction error curves, the corresponding elbow points estimate $m^{\ast}=2$ for all available trajectory data associated with the recurrence phenomenon ($\beta\lesssim 1$) (see Fig. 5).
@@
Finally, the DAE predicts that the intrinsic dimensionality increases by one, i.e., $m^{\ast}=3$, when $\beta=1.1$. We find that this change is associated with the onset of a symmetry-breaking (SB) phenomenon characterizing the $\beta$ model, in which additional energy modes with even wave numbers $k=2,4$ begin to share the total energy of the system. By contrast, this significant variation of ID is completely missed by PCA, for which the estimated dimensionality remains constant as $\beta\to 1.1$.
@@ II FPUT $\beta$ Model
Accordingly, the equations of motion of the $\beta$ model are the following  [^1] The equations of motion are commonly referred to as the canonical Hamilton equations.
@@
From a mathematical point of view, the Cauchy problem for the first-order system in Eq. 2 is well posed; that is, the solution exists and is unique because the vector field is analytic, being polynomial, since the number of oscillators $N$ is finite [58]. However, the solution can be obtained only through numerical integration. To this end, we used the velocity Störmer-Verlet scheme, which is a second-order symplectic integrator well suited for Hamiltonian systems such as FPUT chains [33, 6]. Consequently, in the present work, the trajectories were generated with high accuracy using a fixed time step of 0.05 [30].
@@ III A Deep Autoencoder Model
The autoencoder is an ANN trained to reconstruct an input vector $\mathbf{x}\in\mathbb{R}^{n}$ at the output layer, i.e., to produce a reconstruction $\mathbf{\tilde{x}}$, by passing $\mathbf{x}$ through a narrower hidden layer referred to as the latent (or bottleneck) layer. In this layer, the latent (or code) vector $\mathbf{z}\in\mathbb{R}^{m}$ ($1\leq m\ll n$) provides a compressed, low-dimensional, and salient representation of $\mathbf{x}$. This representation is achieved by minimizing the reconstruction error $J_{m}$ (see Eq.  8), or equivalently, the mean squared error (MSE) loss. In general, AEs can be regarded as nonlinear generalizations of PCA: indeed, it can be shown that an autoencoder restricted to be linear is equivalent to PCA [51, 69].
In the present work, we shall employ a deep autoencoder model with five hidden layers $N_{h}=5$, whose architecture with layer widths $[64,32,16,m,16,32,64]$ is schematically depicted in the cartoon shown in Fig. 2. Note that its architecture consists of two sequential networks: the encoder and the decoder. The first maps $\mathbf{x}$ to $\mathbf{z}$, while the second maps $\mathbf{z}$ to $\mathbf{\tilde{x}}$.
Figure 2: Schematic view of a DAE model with $N_{h}=5$ hidden layers. The input and output layers contain $n$ neurons, while the bottleneck layer has $m$ neurons ($m\ll n$). All hidden neurons use ReLU activation function, as indicated by the symbol. The output layer’s neurons employ a linear activation function, which is appropriate for regression.
In the following, we describe all the model hyperparameters, which were chosen based on their excellent reconstruction performance of the input data for latent dimensions $m\geq 2$, together with a preference for simpler models (Ockham’s razor) [40, 29].
The number of neurons (or nodes) per layer, together with the respective activation functions are shown in Table 1. Note that all hidden neurons use ReLU (Rectified Linear Unit) activation function, i.e., $h(y)=\max(0,y)$, for learning nonlinear mappings [62]. By contrast, the output layer employs a linear activation function, which is commonly used for regression tasks.
Layer Type Number of Neurons Activation Function
@@
bottleneck layer $m$ ReLU
@@
Table 1: Number of neurons and corresponding activation functions of the DAE architecture under consideration. The size of the bottleneck layer, $m$, is assumed to satisfy $1\leq m\ll n$, with $n=64$.
@@
The model is trained using the Adam optimizer [44] with a learning rate of $0.001$ and a mini-batch size of $256$, noting that smaller batch sizes (e.g., $32$) do not produce significant differences in performance. The number of training iterations is set to $300$ epochs. However, early stopping (ES), which acts as an implicit regularizer with a patience of $30$ epochs, is chiefly used to avoid unnecessary training, thereby saving significant time and computational resources. Additionally, the ES will retain the model corresponding to the minimum validation error during the optimization of the trainable parameters (weights and biases)  [^2] Note that when the model’s bottleneck has a single neuron, i.e., $m=1$, there are $4,801$ trainable parameters.. We implemented this model using Keras and TensorFlow [28, 12, 2].
@@
In general, especially for neural networks, the training reconstruction error can be very low because the model may fit the training data extremely closely, thereby leading to serious overfitting. As a result, the model may not perform well on unseen data. Accordingly, to ensure that the model generalizes successfully, we split each trajectory dataset into a training set and a test set [35]. The training set contains $70\%$ of the input data, consisting of data points chosen at random, while the test set contains the remaining $30\%$ of the data. Note that $20\%$ of the training set is also used as validation set for monitoring the model’s generalization during training. With this choice, we find that the difference between the training and test reconstruction errors is generally much less $1\%$.
To estimate ID from the reconstruction error curves [10], we need to restrict the latent dimensionality to as low as a single variable ($m=1$). This limitation is inherent to the constraint $m=1$ itself, and holds independently of the specific autoencoder architecture employed as discussed in next Sec. IV. In such a case, it would be unrealistic to expect an autoencoder to capture the complex structure of the $64$-dimensional input data, unless the data lie on an approximately one-dimensional manifold [51], which is not necessarily the case since we estimated ID to be greater than unity ($m^{\ast}=2,3$). Accordingly, this setup will cause a severe information-bottleneck regime, in which the DAE model is expected to perform poorly, but still significantly better than PCA. However, the aforementioned limitation disappears altogether when the latent dimension $m\geq 2$.
It is worth noting that there is considerable room for improvement over the present model, for example by employing learning-rate decay, alternative activation functions such as GeLU, Mish, or SeLU, or a deeper network architecture [24]. To illustrate this, a comparison with two different deeper models with $N_{h}=7$ is briefly addressed in the next section (Sec. IV). Nevertheless, our preference for a simpler model is justified by the reasonably accurate results achieved at a relatively low computational cost.
@@ III.1 Datasets and Data Preprocessing
Each dataset of size $n_{s}=4\,000\,000$ corresponds to a set $\mathcal{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{i},\cdots,\mathbf{x}_{n_{s}}\}$, whose elements $\mathbf{x}_{i}\in\mathbb{R}^{n}$ are data points that belong to a single trajectory obtained from numerical simulations of the FPUT system at a given value of $\beta\in[0.1,1.1]$, sampled at intervals of $\Delta\beta=0.1$ [57].
The set $\mathcal{X}$ is then conveniently arranged in a $n_{s}\times n$ data matrix $X$. However, before feeding it into the DAE model, a data preprocessing step is required, consisting of centering and scaling each variable by its standard deviation, as is customary in PCA [42]. First, centering the data, corresponding to setting the mean of each variable to zero, allows the PCA eigenvalues to be efficiently computed via singular value decomposition (SVD) (see Sec. B). Second, rescaling ensures that each variable, now having unit variance, contributes equally to the analysis regardless of its original scale [42]. This preprocessing step allows for a direct comparison between the DAE results and those obtained from the principal component analysis in Ref. [57].
@@ IV Results and Discussion
In the following, we first compare the performance of different DAEs to demonstrate that the choice of the model described in Section III is reasonable. Second, we focus specifically on the performance of our model. To this end, we study the behaviour of the learning curves obtained during model training on a specific trajectory dataset corresponding to $\beta=0.3$. However, this analysis exemplifies the characteristic behaviour of the model across all the datasets under consideration.
To begin, we compare our relatively simple model with two others featuring a deeper network structure ($N_{h}=7$), with layer widths given by $[64,48,32,16,m,16,32,48,64]$ for bottleneck sizes $m=1$ and $m=2$. The first model differs only in the value of $N_{h}$, while the second introduces additional features: the GeLU (Gaussian Error Linear Unit) activation function and the SOAP (Simultaneous Orthogonalization and Preconditioning) optimizer with an exponential learning rate, replacing ReLU and Adam, respectively. The SOAP optimizer is a recently proposed method that accounts for correlations between parameters, whereas Adam treats them as independent [79]. The DAE incorporating SOAP, GeLU, and an exponential learning rate was suggested and implemented by Froheberg using the JAX library [8] [^3] The learning rate follows an exponential decay schedule with an initial value of $10^{-3}$, decay rate $0.9$, and transition steps of $10^{5}$..
@@
Figure 3: Training and validation MSE loss curves as functions of epoch for bottleneck sizes $m=1,2$ and trajectory data with $\beta=0.3$ from DAE models with (ReLU, $N_{h}=5$; widths $[64,32,16,m,16,32,64]$) and (ReLU, $N_{h}=7$; widths $[64,48,32,16,m,16,32,48,64]$) using the Adam optimizer, and a DAE model with (GeLU, $N_{h}=7$; widths $[64,48,32,16,m,16,32,48,64]$) using the SOAP optimizer. Throughout training, all deep autoencoders employ early stopping with a patience of $30$ epochs.
Figure 4: Training and validation MSE loss curves as functions of epoch for bottleneck sizes $m=1,2$ and trajectory data with $\beta=0.3$ from DAE model with (ReLU, $N_{h}=5$, $[64,32,16,m,16,32,64]$) using Adam optimizer. (Inset) Validation MSE loss curves as functions of epoch for $m=3,4,5$ and same data. Throughout training, the deep autoencoder employs an early stopping with a patience of $30$.
In Fig. 4 the training and validation MSE losses from the different DAEs are shown as functions of epoch for bottleneck sizes $m=1$ and $m=2$. Throughout, ES is set to $30$. The learning curves display significant qualitative and quantitative differences according to the latent dimension under scrutiny. Overall, we do not observe overfitting, as validation curves follow the training curves closely, yielding nearly identical errors. Accordingly, the models generalize well to unseen data.
When $m=1$, regardless of the chosen model, the irregular oscillations observed after a fast initial decay suggest that these autoencoders are approaching their capacity limit, signalling their inability to faithfully reconstruct the high-dimensional data. Overall, they perform similarly, yielding a best MSE of approximately 0.0360.0360.0380.038.
In contrast, when $m=2$, all curves display the expected learning trend: after a fast initial decay they plateau. In this case, the model with GeLU activation outperforms the other two, achieving a best MSE $\approx 4\times 10^{-4}$, roughly one order of magnitude lower than the ReLU-based models ($\approx 45\times 10^{-3}$). However, the learning curves of our model ($N_{h}=5$) are smoother than those of the deeper model ($N_{h}=7$), suggesting that increasing the network depth alone does not guarantee better performance without further architectural or training adjustments.
While all models’ errors are extremely low, especially when compared to those from PCA (see discussion below), the fact that the errors for $m=2$ can be as much as $80$ times smaller than those corresponding to $m=1$ strongly suggests that the intrinsic dimensionality exceeds one.
Next, in Fig. 4 the training and validation MSE losses from our chosen DAE with $N_{h}=5$ and ReLU, are shown as functions of epoch for bottleneck sizes $m=1,2,\cdots,5$. In the main panel, it is now possible to better visualize the curves for $m=1$ and $2$. In both cases, the respective curves exhibit a rapid initial decay followed by plateau, indicating that the model quickly learns the dominant structure of the data before reaching an error floor. Moreover, for each value of $m$, the training and validation curves remain close throughout training, suggesting that the DAE generalizes.
However, when the bottleneck dimension is reduced to $m=1$, the loss saturates at a significantly higher value than for $m=2$. Specifically, considering that early stopping is triggered at around epoch $116$, one finds that the minimum validation MSE loss $\approx 0.036$ is attained at epoch $86$. This is roughly $7$ times larger than the corresponding best MSE loss observed when $m=2$  [^4] The best validation MSE loss is $\approx 0.0049$, attained at epoch $101$ when $m=2$.. Additionally, the model’s capacity limitation due to the severe scalar information bottleneck ($m=1$) is further evidenced by the noisier evolution of the loss curve, including the presence of spikes and late-epoch fluctuations. These observations are indicative of optimization difficulties induced by this regime and cannot be substantially mitigated without increasing the latent dimensionality. By contrast, for $m\geq 2$ the model performs well, yielding substantially smaller reconstruction errors, as further supported by the well-converged validation loss curves for $m=3,4,5$ shown in the inset of Fig. 4.
@@
Next, the DAE reconstruction errors $J_{m}$ as a function of the latent dimension $m$ for a given value of $\beta$, form a curve whose shape can help identifying the trajectory data’s intrinsic dimensionality. Accordingly, the ID can then be inferred from the elbow point of such a curve [10]. This heuristic is based on the assumption that, beyond the elbow, $J_{m}$ no longer decreases significantly as $m$ increases [77]. Note that this approach is inherently subject to a degree of ambiguity and subjectivity, but is still routinely used in the machine learning field, e.g., for manifold reduction and clustering [28]. In the following, we shall automate the identification of the elbow using the Kneedle algorithm (KA) [73], although in the present case the elbow points can be easily identified by visual inspection (which is not the case for strong nonlinearities; see discussion below).
@@
Figure 5: (Top panel) The DAE test reconstruction error curves as functions of the dimension $m$ for data corresponding to $\beta=0.1,0.4,0.7,0.8,0.9,1.0$, and $1.1$. The PCA reconstruction error curve ($\beta=1.1$) is rescaled by a factor of $100$ for proper visualization. (Bottom panel) The DAE reconstruction error curve for data corresponding to $\beta=0.1$.